Market value ratios

Ratios are mostly calculated using data drawn from the financial statements of a firm. However, another group of ratios, called market value ratios, relate to a firm’s observable market value, stock prices, and book values, integrating information from both the market and the firm’s financial statements.

Consider the case of Cold Goose Metal Works Inc.:

Cold Goose Metal Works Inc. just reported earnings after tax (also called net income) of $8,000,000 and a current stock price of $28.50 per share. The company is forecasting an increase of 25% for its after-tax income next year, but it also expects it will have to issue 1,500,000 new shares of stock (raising its shares outstanding from 5,500,000 to 7,000,000).

If Cold Goose’s forecast turns out to be correct and its price/earnings (P/E) ratio does not change, what does the company’s management expect its stock price to be one year from now? (Round any P/E ratio calculation to four decimal places.)

$28.11 per share

$28.50 per share

$21.08 per share

$35.14 per share

One year later, Cold Goose’s shares are trading at $52.08 per share, and the company reports the value of its total common equity as $39,354,000. Given this information, Cold Goose’s market-to-book (M/B) ratio is   .

Is it possible for a company to exhibit a negative EPS and thus a negative P/E ratio?

No

Yes

Which of the following statements is true about market value ratios?

Low P/E ratios could mean that the company has a great deal of uncertainty in its future earnings.

High P/E ratios could mean that the company has a great deal of uncertainty in its future earnings.
Answer #1

Current earnings per share = net income/number of shares outstanding

= 8,000,000/5,500,000

=$1.4545

Market price per share =$28.50

P/E ratio = market price per share/earnings per share

=28.50/1.4545

=19.5938

Expected earnings per share =8,000,000*125%/7,000,000

=$1.4286

Market price per share = 1.4286*19.5938

=$27.9917

I.e. $28.11 per share approx.

Book value per share = 39,354,000/7,000,000 =$5.622

Market to book value ratio = 52.08/5.622

=9.2636 times

Yes

Because stock.price can never be negative

Low PE ratios could mean that the company has a great deal of uncertainty in its future earnings